# Distributional Scaling Laws for Emergent Capabilities

### v1

@@ 2.2 Natural Language Tasks
We conduct continued pretraining on LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024). We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
@@ 3 Synthetic Experiment Results
As noted earlier, existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the reverse order addition task from Section 2.1.
@@ 3.3 Bimodality Emerges Abruptly
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth increases from 2 to 3 layers or when model width increases from 1 head to 2 heads. This suggests that our Wasserstein distance approach can identify the moment when a new capability is unlocked. At this moment, the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point where the mode in Figure 3 (top left) increases sharply, which always occurs. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
@@ 3.4 Count: An Instance of U-shaped Scaling
Figure 6: As in Figure 3, we capture various summary statistics for test length 60 on the count task. For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case. In the trend for fixed depth (scaling width), we see a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
@@ 3.5 Is Bimodality a Mirage?
In Figure 8, we plot a histogram for this continuous metric averaged across all samples. The distribution of this metric across random seeds is often still bimodal, evidence that bimodality is not due to thresholding alone. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
@@
Figure 9: Violin plots of last-layer reinitialized Qwen2.5 models on the MMLU benchmark. Left Models are trained on different mixtures of the C4 news subset and the MMLU auxiliary training set. With insufficient MMLU training data, Qwen2.5-0.5B fails to follow the multiple-choice format, resulting in trivial performance. As the MMLU data ratio increases, models trained on different seeds exhibit a bimodal distribution in performance. With enough MMLU examples, all models score above the random baseline. Right: For a fixed data mix, the smaller model (Qwen2.5-0.5B) is constrained by its capacity and shows a bimodal distribution in MMLU performance. In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding. However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the full MMLU capability of the original model.
@@ Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. RZ, TQ, and SK acknowledge support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. RZ is supported by a Kempner Institute Graduate Research Fellowship, Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. TQ and DAM are partially supported by the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship.

### v2

@@ 2.2 Natural Language Tasks
We train LLMs to answer multiple-choice questions, an established breakthrough capability (Srivastava et al., 2023; Snell et al., 2024). Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds.
@@ 3 Synthetic Experiment Results
Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales (Wei et al., 2022), implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the reverse order addition task from Section 2.1.
@@ 3.3 Bimodality Emerges Abruptly
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
@@ 3.4 Count: An Instance of U-shaped Scaling
Figure 6: As in Figure 3, we capture various summary statistics for test length 60 on the count task. For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case. The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
@@ 3.5 Is Bimodality a Mirage?
In Figure 8, we plot a histogram for this continuous metric averaged across all samples. The distribution of this metric across random seeds is often still clustered; its bimodality is not due to thresholding alone. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
@@
Figure 9: Violin plots of last-layer reinitialized Qwen2.5 models on the MMLU benchmark. Left Models are trained on different mixtures of the C4 news subset and the MMLU auxiliary training set. With insufficient MMLU training data, Qwen2.5-0.5B fails to follow the multiple-choice format, resulting in trivial performance. As the MMLU data ratio increases, models trained on different seeds exhibit a bimodal distribution in performance. With enough MMLU examples, all models score above the random baseline. Right: For a fixed data mix, the smaller model (Qwen2.5-0.5B) is constrained by its capacity and shows a bimodal distribution in MMLU performance. In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding. However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the multiple choice capability of the original model.
@@ Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. RZ, TQ, and SK acknowledge support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. RZ is supported by a Kempner Institute Graduate Research Fellowship, Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. TQ and DAM are partially supported by the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship. Our work has been improved by invaluable discussion with Will Merrill and David Chiang.